# Describing Globally Distributed Software Architectures for Tax Compliance

### 01 Dec 2023 v1

@@ IV-A International taxation standards as concern and tax authorities as stakeholder
Figure 2: A schematic overview of a globally distributed software architecture developed by the fictional devnullsoft Group, a multinational software enterprise with three subsidiaries developing software. Using code components across different jurisdictions is taxable.
Only the German devnullsoft GmbH sells this resulting product to customers. Without any further consideration, solely the German devnullsoft GmbH generates profits, which are then fully taxed in Germany according to German law. The Swedish and English tax authorities are left out in the cold because devnullsoft AB and devnullsoft Ltd. have no share of the profit that could be taxed in Sweden or in the UK, although both subsidiaries maintain, test, and develop, or simply own components that are crucial for the software product that made the success of the software possible.
To avoid this scenario and to provide a common ground for international taxation, reducing uncertainty for multinational enterprises, and preventing tax avoidance through profit shifting, nearly all countries in the world agreed on and implemented the so-called arm’s length principle as defined in the OECD Transfer Pricing Guidelines for Multinational Enterprises and Tax Administrations [2]. This has become the guiding principle and the de-facto standard for the taxation of multinational enterprises that requires associated enterprises to operate as if not associated and regular participants in the market from a taxation perspective. To comply with the arm’s length principle, multinational enterprises must pay a so-called transfer price. Transfer prices are the prices at which an enterprise transfers physical goods or—for software much more relevant—intangibles or provides services to associated enterprises. Those transfer prices are established on a market value basis and aim to avoid profit shifts from high to low tax regions. Although other types of intangibles, like cross-border code contributions, code reviews, bug reports, etc., are also relevant from a taxation point of view, we focus in this article on another highly relevant type of intangible in software engineering: licenses.
@@
In this way, tax authorities become a stakeholder in globally distributed software architectures as they have the concern of identifying cross-border reuse of software components as an instance of taxable licensing. Determining the actual transfer pricing for the cross-border reuse of software components is a vast and separate challenge on its own and not part of this work. Our work focuses on the architectural description comprising all necessary information for reporting the software architecture to tax authorities.
@@ IV-B Architecture viewpoint
IV-B1 Defining component and component owner
@@
Component ownership refers to the concept of assigning responsibility and accountability for a particular software component to an individual or an organizational unit within an organization. Although there are different notions of code ownership (for example, strong, weak, or collective code ownership [17]), component ownership is more strictly defined in this context: We requires only one organizational unit is ultimately responsible and accountable for a software component.
@@
IV-B2 Decomposing the component structure of the software product
To reveal the component structure and the dependencies of software components, the software product must be decomposed into the component structure. This results in a dependency graph reflecting the dependencies between the software components. The type of dependency depends on the definition of a software component. Therefore, a rationale is required.
IV-B3 Identifying component owners
@@
IV-B4 Identifying jurisdictions of component owners
@@ IV-C Architecture view
In detail, we use microservices as the most suitable abstraction layer for software components for the specific software architecture in our case. A microservice represents a discrete piece of a larger application’s functionality and operates as a self-contained unit. Microservices are typically designed to be lightweight, modular, and independently scalable, allowing for greater flexibility and agility in software development [15]. Therefore, microservices meet the definition of a software component.
@@
In our case company, each microservice is usually owned by a single team. We excluded 42 microservices (1.64%), each owned by one individual because the locations of individuals were not accessible for this research due to privacy concerns. Those remaining 2518 microservices have 16533 dependencies between them and are owned by 336 different teams.
@@
Figure 3: The geographical distribution of the microservice architecture of our case company using our architectural viewpoint. The arrows indicate the direction of the dependency from microservice user to the microservice owner. The countries are partially substituted by equivalent countries to ensure confidentiality for the company.
This graph, however, lacks of the information of unclear locations. Therefore, we provide also a tabular form of the graph including the information of missing or unclear location. Table II depicts this information.
TABLE II: An equivalent tabular representation of Figure 3 but including unknown dependencies.
@@
V Evaluation results
In this section, we report the evaluation results from the panel of tax experts (see Section III for more details) evaluating the software architecture description with its viewpoint that framed the concerns of tax authorities and the view of software architecture from our industry partner.
V-A Viewpoint
Our panel confirmed the correctness of our viewpoint and agreed with the distilled three questions raised from the concern of international taxation. As interviewee A stated: I think it is possible to reveal the license structure [by your viewpoint], at least by 90%, there will always be fuzziness.
@@ V-A Viewpoint
Our panel emphasized that although aggregating over jurisdictions can be acceptable in our case since there is only one subsidiary per jurisdiction, it does not generalize: If there is more than one subsidiary in a jurisdiction, aggregating over countries is not applicable. Interviewee K also mentioned that assigning legal entities to jurisdiction can become complicated referring to so-called permanent establishments, where, in complex situations, subsidiaries may have international branches and sites fall into jurisdiction X but belong to jurisdiction Y. Assigning jurisdiction is complex. There are gray areas. However, those special cases are likely not relevant for those considerations.
@@ V-B View
The first major issue refers to the problem that we were not able to identify the jurisdiction of 41.16% of all owners in our view on the given software architecture. Although the data source, the company-internal platform for teams, is promising (I go along with that one. […] That is, in my opinion, an excellent data source) since the data is not biased from a taxation perspective (interviewee M), the amount of unclear jurisdiction of owners is too high. The experts suggested different additions (jurisdiction of the team manager) or complements (a cascade of different other data sources, for example, a questionnaire sent out to all owners or geographical distribution of the team to approximate the jurisdiction of the owners). Although missing some percentages is no problem, but half of the dependencies missing is too much as interviewee M explicitly said. 50% not assignable implies that 50% of all transfer prices are not properly assigned. […] Without data, we are in vacuo. (interviewee A).
In this context, interviewee M and interviewee A emphasized the importance of traceability in tax audits. Therefore, it would also be better to reveal the process and characteristics of how teams came to their self-assignment. Indications like regular socializing in a certain country can be such an indicator. In this context, the panel agreed that the physical location of the team members and developers can be secondary from a transfer pricing perspective.
@@
Besides those two major issues, the panel also raised further issues. Interviewee R wondered about the effort required to generate those insights. The interviewee emphasized that the cost-benefit ratio must be positive. Another important remark was related to our decision to exclude non-production components. Interviewee M reminded us that our ultimate goal is to identify the value of software components. Although not in non-production or completely experimental, those components can have a significant and not negligible value for the company. Therefore, excluding them removes explanatory power from the software architecture description.
V-C Software architecture description
@@
The overall impression was positive and, according to our panel, the software architecture description fulfilled its main purpose at its current state to a large extent. Although transfer pricing is not a science, it is art, interviewee R correctly summarized our intention as follows: This is just fact-driven, it is no interpretation. When you ask people, they may or may not know who is involved. The nice thing about it that it is not a subjective view, it is about facts, it is about technical details. This is what is used if you like it or not. That is relevant because you have objectivized the contributions [to the software product]. Those are the parties involved. This allows you to have a very objective starting point to consider the potential parties you have to consider in greater detail. The visualization raises awareness of parties involved in the software development. The panel also confirmed the importance of software architecture for taxation: I think software architecture is the right perspective to understand how software is composed conceptionally (interviewee K). Interviewee R highlighted the need for software that can manage the complexity and support the transfer pricing processes. Interviewee M also appreciated the company’s efforts in collecting and presenting the data in a tax audit: A company with such detailed and thorough documentation would be among the top 10 or top 5 percent of what a tax auditor usually sees. Interviewee K agrees: I really like the approach aiming for objectiveness. Let’s face it: Sometimes there are preposterous debates in tax audits.
VI Discussion
@@
We found ambiguous assignments of owners to jurisdiction in about half of all dependencies. We expected that this significant amount of unclear jurisdictions per team create a strong headwind. Statistical estimators cannot be applied to make any assumptions on the jurisdictions because we have a strong bias: The owners with missing or unclear jurisdictions come mostly from teams that have team members from different subsidiaries or are completely virtual. Those are inherently relevant for taxation purposes and need further investigation.
The discussion on ownership, however, cast doubt on our definition of ownership. We learned that the term ownership triggers a certain association in the context of taxation: Tax experts think of legal ownership when they read ownership. Although we emphasized in our viewpoint that we require an owner to be accountable and responsible, the term ownership led to different misunderstandings and triggered in-depth legal discussions that go beyond the scope of this paper. We conclude that our definition of ownership can be indeed too weak: Our understanding of accountability may fall short and does not suffice the legal accountability. Future research needs to address the legal aspects of code (component) ownership.
@@
We are not aware of studies addressing at what rate or to what extent microservices architectures or software architecture in general change over time (a fiscal year), which would also may be not generalizable and context-specific anyway. Since the effort to create the views once the data sources are identified is reasonable, we suggest regularly generating snapshots of the software architecture. Furthermore, we hypothesize that changes in software architecture also may trigger changes in ownership, directly or indirectly.
VII Limitations
In this section, we critically reflect on the two components of our simulation-based approach: the software architecture description from our industry partner and the evaluation of the software architecture description in the contrived setting.
VII-A Software architecture description
We created the view as described in the paper. We relied on historical information extracted from an internal system that extracts the software product’s component structure (which also comprises microservices) from the code repositories. We believe that this system is a reliable and trustworthy data source for extracting the data we need since the data is also used internally for other purposes. The internal team responsible for maintaining the system helped us to avoid misinterpretation of the data. The date of extraction was randomly selected from the first two quarters in 2023. Although we did not collect this data, we assume that the microservice architecture does not change at a daily rate but is stable at least for over weeks.
VII-B Contrived setting
The evaluation of the software architecture description for tax compliance highly depends on our tax panel. The contrived setting and the strict anonymization of the case company required us to remove context and potentially relevant information, which includes the specific business model, the revenue stream, and organizational structure of our case company. Not being able to provide that information limits the generalizability of our evaluation. However, we believe that we provided enough context and information for the panel to provide solid and high-level feedback for the early stages of our research.
@@ VII-B Contrived setting
To maintain strict anonymity for the company, we decided not to include participants from the company. The author team spoke on their behalf. In future work addressing the subsequent questions of valuing software components, the involvement of internals will become crucial. We believe that in this early stage of the software architecture description, the author team who extracted the data brought enough context to the table.
@@
VIII Conclusion
In this study, we proposed and evaluated a novel software architecture description for reporting globally distributed software architecture for tax compliance. The architecture description of our case company reveals that almost 70% of all reuse relationships between the 2560 microservices in our case are cross-border and, therefore, are taxable.
The panel of tax experts found that our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. The panel also appreciated the evidence-based, empirical approach for visualizing the software architectures to tax experts as software-engineering laymen.
However, we identified several challenges that must be addressed in future work. The most critical and likely most complex issue to solve is our insufficient definition and understanding of ownership, which needs to be rethought for taxation purposes in this line of research. Also the question what does or will constitute to the software product and what not deserves further attention. Last, but not least, the subsequent step, adding value to each component, is yet another unsolved problem in the context of taxing software-intensive products. We believe that the software architecture research community can play a vivid role in addressing those challenges and paving the way for tax compliance in software engineering.
@@ VIII Conclusion
At this point, we would like to emphasize that the quality and rigor of the data curation at our case company is extraordinary and, from our experience, not industry standard. Not always, companies can rely on such a solid foundation grounded in historical and accurate data on their software architectures. This uncertainty introduces further risks and lets tax audits stray away from an empirical-based, data-driven discussion on the value of software. Although this research is only the beginning, we recommend companies start continuously collecting information about their software architecture and ownership. Accurately collecting this information is never trivial but nearly impossible in hindsight for complex software systems.
Although there is no black or white in tax audits but a lot of gray, as interviewee R summed up, we hope to shed some light on the discussions in tax audits and lay the foundation for a data-driven approach for taxing globally distributed software architectures.
Acknowledgment

### 04 Mar 2024 v2

@@ 4.1. International taxation standards as concern and tax authorities as stakeholder
Figure 2. A schematic overview of a globally distributed software architecture developed by the fictional devnullsoft Group, a multinational software enterprise with three subsidiaries developing software. Using code components across different jurisdictions is taxable.
Only the German devnullsoft GmbH sells this resulting product to customers. Without any further consideration, solely the German devnullsoft GmbH generates profits, which are then fully taxed in Germany according to German law. The Swedish and English tax authorities are left out in the cold because devnullsoft AB and devnullsoft Ltd. have no share of the profit that could be taxed in Sweden or the UK, although both subsidiaries maintain, test, and develop, or simply own components that are crucial for the software product that made the success of the software possible.
To avoid this scenario and to provide a common ground for international taxation, reducing uncertainty for multinational enterprises, and preventing tax avoidance through profit shifting, nearly all countries in the world agreed on and implemented the so-called arm’s length principle as defined in the OECD Transfer Pricing Guidelines for Multinational Enterprises and Tax Administrations (OECD, 2022). This has become the guiding principle and the de-facto standard for the taxation of multinational enterprises that requires associated enterprises to operate as if not associated and regular participants in the market from a taxation perspective. To comply with the arm’s length principle, multinational enterprises must pay a so-called transfer price. Transfer prices are the prices at which an enterprise transfers physical goods or—for software much more relevant—intangibles or provides services to associated enterprises. Those transfer prices are established on a market value basis and aim to avoid profit shifts from high to low tax regions. Although other types of intangibles, like cross-border code contributions, code reviews, bug reports, etc., are also relevant from a taxation point of view, we focus in this article on another highly relevant type of intangible in software engineering: licenses.
@@
In this way, tax authorities become stakeholders in globally distributed software architectures as they have the concern of identifying cross-border reuse of software components as an instance of taxable licensing. Determining the actual transfer pricing for the cross-border reuse of software components is a vast and separate challenge on its own and not part of this work. Our work focuses on the architectural description comprising all necessary information for reporting the software architecture to tax authorities.
@@ 4.2. Architecture viewpoint
4.2.1. Defining component and component owner
@@
Component ownership refers to the concept of assigning responsibility and accountability for a particular software component to an individual or an organizational unit within an organization. Although there are different notions of code ownership (for example, strong, weak, or collective code ownership (Fowler, 2006)), component ownership is more strictly defined in this context: We require only one organizational unit to be ultimately responsible and accountable for a software component.
@@
4.2.2. Decomposing the component structure of the software product
The software product must be decomposed into the component structure to reveal the component structure and the dependencies of software components. This results in a dependency graph reflecting the dependencies between the software components. The type of dependency depends on the definition of a software component. Therefore, a rationale is required.
4.2.3. Identifying component owners
@@
4.2.4. Identifying jurisdictions of component owners
@@ 4.3. Architecture view
In detail, we use microservices as the most suitable abstraction layer for software components for the specific software architecture in our case. A microservice represents a discrete piece of a larger application’s functionality and operates as a self-contained unit. Microservices are typically designed to be lightweight, modular, and independently scalable, allowing for greater flexibility and agility in software development (Newman, 2021). Therefore, microservices meet the definition of a software component.
@@
In our case company, each microservice is usually owned by a single team. We excluded 42 microservices (1.64%), each owned by one individual, because the locations of individuals were not accessible for this research due to privacy concerns. Those remaining 2518 microservices have 16533 dependencies between them and are owned by 336 different teams.
@@
Figure 3. The geographical distribution of the microservice architecture of our case company using our architectural viewpoint. The arrows indicate the direction of the dependency from the microservice user to the microservice owner. The countries are partially substituted by equivalent countries to ensure confidentiality for the company.
This graph, however, lacks of the information of unclear locations. Therefore, we also provide a tabular form of the graph, including the information on missing or unclear locations. Table 2 depicts this information.
Table 2. An equivalent tabular representation of Figure 3 but including unknown dependencies.
@@
5. Evaluation results
In this section, we report the evaluation results from the panel of tax experts (see Section 3 for more details) evaluating the software architecture description with its viewpoint that framed the concerns of tax authorities and the view of software architecture from our industry partner.
5.1. Viewpoint
Our panel confirmed the correctness of our viewpoint and agreed with the three questions distilled from the concern of international taxation. As interviewee A stated: I think it is possible to reveal the license structure [by your viewpoint], at least by 90%, there will always be fuzziness.
@@ 5.1. Viewpoint
Our panel emphasized that although aggregating over jurisdictions can be acceptable in our case since there is only one subsidiary per jurisdiction, it does not generalize: If there is more than one subsidiary in a jurisdiction, aggregating over countries is not applicable. Interviewee K also mentioned that assigning legal entities to jurisdiction can become complicated referring to so-called permanent establishments, where, in complex situations, subsidiaries may have international branches and sites fall into jurisdiction X but belong to jurisdiction Y. Assigning jurisdiction is complex. There are gray areas. However, those special cases are likely not relevant to those considerations.
@@ 5.2. View
The first major issue refers to the problem that we were not able to identify the jurisdiction of 41.16% of all owners in our view on the given software architecture. Although the data source, the company-internal platform for teams, is promising (I go along with that one. […] That is, in my opinion, an excellent data source) since the data is not biased from a taxation perspective (interviewee M), the amount of unclear jurisdiction of owners is too high. The experts suggested different additions (jurisdiction of the team manager) or complements (a cascade of different other data sources, for example, a questionnaire sent out to all owners or geographical distribution of the team to approximate the jurisdiction of the owners). Although missing some percentages is no problem, but half of the dependencies missing is too much as interviewee M explicitly said. 50% not assignable implies that 50% of all transfer prices are not properly assigned. […] Without data, we are in a vacuum. (interviewee A).
In this context, interviewee M and interviewee A emphasized the importance of traceability in tax audits. Therefore, it would also be better to reveal the process and characteristics of how teams came to their self-assignment. Indications like regular socializing in a particular country can be such an indicator. In this context, the panel agreed that the physical location of the team members and developers can be secondary from a transfer pricing perspective.
@@
Besides those two major issues, the panel also raised further issues. Interviewee R wondered about the effort required to generate those insights. The interviewee emphasized that the cost-benefit ratio must be positive. Another critical remark was related to our decision to exclude non-production components. Interviewee M reminded us that our ultimate goal is to identify the value of software components. Although not in non-production or completely experimental, those components can have a significant and not negligible value for the company. Therefore, excluding them removes explanatory power from the software architecture description.
5.3. Software architecture description
@@
The overall impression was positive and, according to our panel, the software architecture description fulfilled its main purpose at its current state to a large extent. Although transfer pricing is not a science, it is art, interviewee R correctly summarized our intention as follows: This is just fact-driven, it is no interpretation. When you ask people, they may or may not know who is involved. The nice thing about it is that it is not a subjective view; it is about facts, it is about technical details. This is what is used, if you like it or not. That is relevant because you have objectivized the contributions [to the software product]. Those are the parties involved. This allows you to have a very objective starting point to consider the potential parties you have to consider in greater detail. The visualization raises awareness of parties involved in the software development. The panel also confirmed the importance of software architecture for taxation: I think software architecture is the right perspective to understand how software is composed conceptionally (interviewee K). Interviewee R highlighted the need for software that can manage the complexity and support the transfer pricing processes. Interviewee M also appreciated the company’s efforts in collecting and presenting the data in a tax audit: A company with such detailed and thorough documentation would be among the top 10 or top 5 percent of what a tax auditor usually sees. Interviewee K agrees: I really like the approach aiming for objectiveness. Let’s face it: Sometimes, there are preposterous debates in tax audits.
6. Discussion
@@
We found ambiguous assignments of owners to jurisdiction in about half of all dependencies. We expected this significant amount of unclear jurisdictions per team to create a strong headwind. Statistical estimators cannot be applied to make any assumptions on the jurisdictions because we have a strong bias: The owners with missing or unclear jurisdictions come primarily from teams with team members from different subsidiaries or are entirely virtual. Those are inherently relevant for taxation purposes and need further investigation.
The discussion on ownership, however, cast doubt on our definition of ownership. We learned that the term ownership triggers a particular association in the context of taxation: Tax experts think of legal ownership when they read ownership. Although we emphasized in our viewpoint that we require an owner to be accountable and responsible, the term ownership led to different misunderstandings. It triggered in-depth legal discussions that go beyond the scope of this paper. We conclude that our definition of ownership can be indeed too weak: Our understanding of accountability may fall short and does not suffice the legal accountability. Future research must address the legal aspects of code (component) ownership.
@@
We are not aware of studies addressing at what rate or to what extent microservices architectures or software architecture in general change over time (a fiscal year), which also may be not generalizable and context-specific anyway. Since the effort to create the views once the data sources are identified is reasonable, we suggest regularly generating snapshots of the software architecture. Furthermore, we hypothesize that changes in software architecture also may trigger changes in ownership, directly or indirectly.
7. Limitations
In this section, we critically reflect on the two components of our simulation-based approach: the software architecture description from our industry partner, the evaluation of the software architecture description in the contrived setting, and the required anonymity.
7.1. Software architecture description
We created the view as described in the paper. We relied on historical information extracted from an internal system that extracts the software product’s component structure (which also comprises microservices) from the code repositories. We believe that this system is a reliable and trustworthy data source for extracting the data we need since the data is also used internally for other purposes. The internal team responsible for maintaining the system helped us to avoid misinterpretation of the data. The date of extraction was randomly selected from the first two quarters in 2023. Although we did not collect this data, we assume that the microservice architecture does not change at a daily rate but is stable, at least for over weeks.
7.2. Contrived setting
The evaluation of the software architecture description for tax compliance highly depends on our tax panel. The contrived setting and the strict anonymization of the case company required us to remove context and potentially relevant information, which includes the specific business model, the revenue stream, and the organizational structure of our case company. Not being able to provide that information limits the generalizability of our evaluation. However, we believe that we provided enough context and information for the panel to provide solid and high-level feedback for the early stages of our research.
@@ 7.2. Contrived setting
7.3. Anonymization
To maintain strict anonymity for the company, we decided not to include participants from the company. The author team spoke on their behalf. We believe that in this early stage of the software architecture description, the author team who extracted the data brought enough context to the table. In future work addressing the subsequent questions of valuing software components, the involvement of internals, however, will become crucial.
We believe that anonymization has a substantial influence on the quality of the evaluation, though. We removed (intentionally) the context of the software architecture. Of course, this context also plays a crucial role in taxation since software engineering does not happen in the void. In particular, the economic context could not be revealed, which led to several loose ends during the evaluation. We believe that this is a major limitation of our findings since the devil is in the details. We could not investigate the potential effects of economic or managerial decisions on the software architecture.
8. Conclusion
We are torn for the conclusion of this research.
@@ 8. Conclusion
On the one hand, our novel software architecture description for reporting globally distributed software architecture for tax compliance reveals that almost 70% of all reuse relationships between the 2560 microservices in our case are cross-border and, therefore, taxable. The extent is magnitudes higher than what we would have expected. Therefore, we urge multinational enterprises to start collecting information about their software architecture and ownership continuously. Accurately collecting this information is never trivial but nearly impossible in hindsight for complex software systems. At this point, we would like to emphasize that the quality and rigor of the data curation at our case company is extraordinary and, from our experience, not industry standard. Not all companies can rely on such a solid foundation grounded in historical and accurate data on their software architectures. This uncertainty introduces further risks and lets tax audits stray away from an empirical-based, data-driven discussion on the value of software. The panel of tax experts found that our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. Although the resulting view of the globally distributed software architecture of our industry partner has to cope with a significant amount of uncertainty, tax experts are used to working with uncertainty (there is no black or white in tax audits but a lot of gray, as interviewee R summed up) and appreciate the evidence-based, empirical approach for visualizing the software architectures to tax experts as software-engineering laymen.
On the other hand, however, we were surprised to be confronted by more fundamental questions that still remain unanswered in software engineering (research). Those questions touch both, currently insufficient definitions of core concepts (such as code ownership or software component), but also limitations in our ways of doing research itself. Although we believe that our research laid the foundation for a data-driven approach to tax compliance of globally distributed software architectures, we believe that tax compliance in software engineering requires us, the community of interested researchers, to take a step backwards first. In the following subsections, we discuss questions we believe are important to address first and the avenue towards this new frontier of taxing software engineering.
8.1. What is code ownership?
The probably most critical and likely most complex issue to solve is our insufficient definition and understanding of code ownership, which needs to be rethought and extended for taxation purposes. As we learned in the course of this study (see Section 5), the definition of ownership in the context of software engineering and the definition of ownership in taxation overlap but are not congruent due to the different contexts and perspectives towards ownership.
Ownership in software engineering, on the one hand, refers to the assignment of responsibility and accountability for specific portions of software to individuals or teams within a development organization. There are different perspectives on ownership and autonomy in software development organizations. Software development organizations usually adopt the weak ownership principle (Fowler, 2006). This principle entails assigning ownership of specific components (or microservices) to individual teams. In an ideal scenario, the designated team assumes responsibility as the primary (or exclusive) contributor to the associated component. Nevertheless, empirical analysis of real-world instances often reveals a departure from this idealized model (Borg et al., 2023; Tanveer et al., 2023; Greiler et al., 2015; Bird et al., 2011). In practice, a given component may offer support to multiple business streams, undergo development by distinct teams, and necessitate the involvement of specialized teams to address non-functional requirements, such as security and reliability. Consequently, while a specific team retains responsibility for the component, the extent of its contribution to the component can exhibit considerable variability (Zabardast et al., 2022).
Thongtanunam et al. argues that code ownership comes with responsibility, and developers who author the majority of changes to a module are presumably the owners (Thongtanunam et al., 2016). However, contribution may come in other forms, e.g., in modern code review by criticizing code changes authored by other developers, and the study was based on open-source projects. We also believe that results found in open-source projects cannot be directly transferred without further consideration due to the differences in governance structures, incentives, and economic mechanics.
Tax professionals, on the other hand, begin from a civil law perspective and understand ownership in terms of the legal or economic owner regarding the right of exploitation (OECD, 2022). While the identification of the legal or economic owner is relevant for balance sheet allocations, in the context of transfer pricing, the legal owner is considered (only) as the starting point for a functional and risk analysis that aims for a profit allocation commensurate with the value-adding activities performed by the legal entities within the group, the so-called functional ownership (OECD, 2022).
We believe that describing and discussing the intersection, as well as the difference in definitions, requires further more interdisciplinary research. This work is not per-se a purely academic exercise but an endeavour with significant practical relevance for tax compliance.
8.2. What is a software component?
Although the definition of software components are vague and context-specific (for example, (Broy et al., 1998)), the hierarchical, tree-like nature of components and, therefore, software architectures, makes us worrying that we the level of abstraction (i.e. microservices in this research) is the optimal choice.
First, it goes without saying, that also non-microservice software architectures are subject to tax compliance.
Second, our architecture description is, like all models, a simplified abstraction of reality. This simplified abstraction enables us to understand the reality or, in our case, derive taxable conclusions from the software architecture. No model fits all purposes and can, therefore, be perfect, and all are prone to error. This model error is a result of the superposition of two error curves: (i) the bias error, which results from too simple models and can be mitigated by increasing the model complexity, and (ii) the variance error, which reduces the uncertainty with growing model complexity (see Figure 4).
[…]
Figure 4. The overall model error is a result from the superposition of two error curves: (i) the bias error, which results from too simple models and can be mitigated by increasing the model complexity, and (ii) the variance error, which reduces the uncertainty with growing model complexity. We are unsure if we hit the optimal model complexity with our software architecture description.
For our work, this means that we might have abstract software architecture on a too-high-level understanding of software components or a too-low level of software components. On the one hand, high-level understandings such as microservices or even systems (which would then lack a clear definition again) leave more—potentially relevant—information unmodelled but leave the model complexity reasonable. For example, we use only one unspecific relationship between components: the use relationship, which neglects the detailed characteristics of a use relationship or the tree-like nature of software components, resulting in more than 2500 microservices. On the other hand, a low-level understanding of software components (for example, on file level) results in very complex and gigantic models describing the software architecture. In such a software architecture description, the inherent uncertainty would be even more salient.
Therefore, we are unsure if our software architecture description is prone to underfitting or overfitting.
8.3. How can we report empirical evidence?
Many research endeavours that emerge from academia-industry collaborations are prone to the challenge that not all data can be shared “as-is” for confidentiality reasons. We would normally argue for, at least, partial disclosure, anonymization of disclosed data, or sharing of metadata (and, where reasonable, making data conditionally available upon request).
For this research line, however, absolute anonymity and confidentiality between the researchers and all involved parties are paramount to foster trust, honesty, and candor among the participants to ensure that individuals freely contribute highly sensitive information without fear of identification or reprisal. While only an absolute dedication to the anonymity of our case and all involved subjects can safeguard confidentiality and privacy, it also poses a natural barrier to any data dissemination and, more generally, to open science, which we ourselves hold dear as the de-facto standard for empirical software engineering (Mendez et al., 2020).
In our line of research, however, we are simply not able to reveal any details about the company, the software architecture, or the interviews without potentially revealing the identity of the involved parties. Even how we collected the data or what data source we used would pose the risk, even if remotely, of revealing the identity of the participating company. This constitutes a fundamental barrier to reproducibility, also in light of that we deem a replication with publicly available data not feasible nor ethical (also considering that we can also not make a case for using open source projects which we discussed already to be not comparable). At the same time, we may well argue that our reporting is still suitable to uncover the potential of describing globally distributed software architectures for the purpose of critically analyzing tax compliance and, at the same time, more fundamental questions that open an interesting avenue towards this new frontier of taxing software engineering. In any case, this general dilemma still remains, in our view, something we deem important to discuss in the research community of empirical software engineering.
Data Availability
Due to the nature of the research, supporting data, except interview guidelines [^2] https://zenodo.org/records/10214942, is not available. Please see our discussion in Section 8.3.
Acknowledgements.